# Color3D: Controllable and Consistent 3D Colorization with Personalized Colorizer

### v1

@@ A.3 More Experiments
Image and video colorizers vs. per-scene colorizer. We evaluate the effectiveness of our per-scene colorizer design by comparing it with two representative baselines: the image colorization model DDColor Kang et al. (2023) and the video colorization method ColorMNet Yang et al. (2024a), as illustrated in Fig. 9. It can be observed that DDColor suffer from severe cross-view color inconsistencies due to its per-image one-to-many color mapping nature. Treating multi-view images as a video sequence and applying video colorization similarly fails to resolve this issue, since the viewpoint changes inherent in 3D capture far exceed the temporal variations commonly seen in video data. In contrast, our personalized colorizer learns a scene-specific, one-to-one color mapping from the reference view and leverages its inherent inductive bias to consistently project colors to corresponding content in novel views and time steps. This design effectively addresses the challenge of color consistency in both static and dynamic 3D scenes.
@@
Method
Language-Guided Colorization
@@
NeRF+2DColorizer 104.85 0.6153 0.225
3DGS+2DColorizer 112.33 0.6148 0.202
Color3D-NeRF 76.68 0.6213 0.065
Color3D (Ours) 68.23 0.6246 0.058
@@
NeRFPlayer+2DColorizer 86.77 0.6148 0.133
4DGS+2DColorizer 89.39 0.6159 0.124
Color3D-NeRFPlayer 63.29 0.6243 0.046
Color3D (Ours) 58.62 0.6271 0.041
Effort of 3D representations. We further investigate how different 3D representations impact colorization performance. As shown in Tab.3, compared to 3DGS, combining NeRFMildenhall et al. (2021) or NeRFPlayer Song et al. (2023) with 2D colorizers yields suboptimal consistency, reflected by an increase of 0.023 in ME. This is mainly because the explicit modeling of 3DGS gives it an inherent resistance and geometric priors, which can naturally alleviate color inconsistency to some extent. In addition, we replace the Lab Gaussian representation in Color3D with NeRF and NeRFPlayer, respectively. The results underscore the effectiveness of our proposed Lab Gaussian representation while also demonstrating that the Color3D framework can be adapted to various 3D reconstruction backbones with reasonable performance.
@@
User study. To further assess the perceptual quality of the generated results, we conduct a comprehensive user study, acknowledging that subjective human perception remains a critical benchmark in the field of colorization. Specifically, we design 5 evaluation questions covering multiple aspects, including color richness, consistency, aesthetic preference, image quality, and the fidelity of alignment between the synthesized views and the provided conditions (text, exemplars). Participants are asked to rate each sample on a scale of 0 to 5 for each aspect. For a fair comparison, we recruit 30 participants, each evaluating 10 sets of comparisons across different methods. Aggregated results are shown in Fig. 12. The naive combination of 3DGS with off-the-shelf 2D colorizers yields reasonably rich colors but suffers from significant inconsistency and reduced visual quality. While ColorNeRF demonstrates better consistency, it often lacks color diversity and visual appeal. In contrast, our method consistently achieves higher user ratings across all aspects, highlighting its advantage in terms of perceptual quality and user preference.
@@
A.4 Limitations and Future Work

### v2

@@ A.5 More Experiments
DL3DV-10-Benchmark Ling et al. (2024) consists of 140 large-scale 360 real-world scenes. All images are processed at a resolution of $960\times 540$ for our experiments.
@@
HyperNeRF Park et al. (2021b) is another dynamic dataset captured using one or two cameras following relatively straightforward camera trajectories, but with more intricate camera motions.
@@
Quantitative comparison with ChromaDistill. Additionally, following ChromaDistill’s experimental setup, we conduct automatic image colorization experiments using BigColor Kim et al. (2022) across four scenes: Cake, Pasta, Buddha, and Leaves. The experimental results are depicted in Tab. 5. It is observed that our method achieves substantially higher consistency than ChromaDistill in both short-term and long-term consistency. Notably, on the Cake scene, our approach improves long-term consistency by 0.023, further demonstrating the superiority of the proposed personalized colorizer in preserving cross-view coherence and validating the effectiveness of our overall framework.
Table 5: Quantitative comparisons with ChromaDistill on cross-view short-term and long-term consistency.
Method Short-Term Consistency $\downarrow$ Long-Term Consistency $\downarrow$
Cake Pasta Buddha Leaves Cake Pasta Buddha Leaves
ChromaDistill (BigColor) 0.019 0.015 0.015 0.008 0.033 0.025 0.023 0.013
Color3D (BigColor) 0.008 0.009 0.009 0.007 0.010 0.018 0.012 0.011
Single view or more? Due to the inherent instability of 2D colorization models, even small perturbations in the input image can lead to noticeably different color predictions for both individual objects and the overall scene style. When two or more key views are used for training the personalized colorizer, these inconsistencies across views introduce conflicting supervision, which often causes the model to converge to color averaging or to produce view-dependent color shifts during 3D colorization. As shown in Fig. 13, the results obtained by training with two views clearly exhibit such inconsistencies, making it difficult for the personalized colorizer to maintain coherent colors across views. In addition, the quantitative comparison in Tab. 6 shows that using a single key view achieves better consistency and avoids both color drifting and color averaging, leading to superior FID and Colorful scores.
Figure 13: Visual comparison of personalized colorizer training with single vs. two key views.
Table 6: Quantitative comparison of using different numbers of key views for automatic colorization on the Mip-NeRF-360 dataset.
[…]
Three Views 52.38 25.26 0.098 0.022
Two Views 45.62 27.36 0.092 0.018
One View 39.03 33.36 0.082 0.016
Robustness to key view selection. Our framework demonstrates strong robustness with respect to the choice of the key view. As illustrated in Fig. 14, varying the selected key view consistently leads to coherent and chromatically vibrant 3D colorization results. Even when the first key view contains almost no grass, our approach can still produce reasonable and vibrant colors. Moreover, for distant buildings that do not appear in the key view at all, our method can still infer plausible colors for them. These can be primarily attributed to the generative data augmentation strategy and the colorizer’s ability to interpolate colors for unseen content, which together enable stable colorization even when the selected key view is not ideal.
Figure 14: Visual comparison of the colorization results with different key views.
What is the color of semantically similar but visually different objects? To further evaluate the generalization capability of our personalized colorizer, we conduct an experiment on scenes containing semantically similar yet visually distinct objects. As shown in Fig. 15, our method consistently assigns plausible and stable colors even when object appearances in the target views deviate from those seen in the key view. This behavior arises from two key design choices. First, the frozen encoder, pre-trained on large-scale datasets, provides robust and viewpoint-invariant visual features. During optimization, the encoder yields consistent representations for the same object across different viewpoints while still capturing feature differences for objects that are semantically similar but visually distinct. Second, the decoder, trained jointly on the key view and our generatively augmented images, learns a broader color mapping space. This enables it to perform effective color interpolation when encountering unseen visual features.
Figure 15: Visual comparison of the colorization results on objects that share semantic similarity but differ in visual appearance. More views can be found in Fig. 23
Performance under challenging lighting conditions. We also demonstrate the performance of our method under challenging lighting conditions in Fig. 16. Specifically, we evaluate scenes with complex lighting, including lamps, reflections, shadows, and highly specular materials. Our results show that the proposed method effectively handles these challenging scenarios and produces consistent and reasonable colorization even in regions affected by shadows and projections. These findings further validate the reliability and effectiveness of our approach in practical applications.
Figure 16: Visual results of our method under challenging lighting conditions.
Optimization efficiency. In Tab. 7, we compare the optimization time across different methods. While our Stage 1 introduces an extra 8 minutes, the Stage 2 optimization is significantly faster than that of 3DGS+ImageColorizer, and the overall time is even shorter than reconstructing a 3D scene from normal colored images using 3DGS. This efficiency is mainly due to our proposed Lab-space optimization strategy, which first optimizes the L channel to capture structural information and then learns color modeling. This strategy substantially reduces redundant Gaussian clones and splits, allowing fewer Gaussian primitives to model the scene more effectively. In contrast, 3DGS+ImageColorizer splits a large number of Gaussian primitives to model multi-view inconsistent colors, resulting in longer optimization times than standard 3DGS reconstruction. ColorNeRF requires an even longer 16 hours to reconstruct a single scene. These results demonstrate that our method is more efficient than baseline reconstruction approaches, and that per-scene optimization is acceptable in practice.
Table 7: Average optimization time on Mip-NeRF-360 (static) and DyNeRF (dynamic) datasets.
Dataset Ours 3D/4DGS ColorNeRF
Stage 1 Stage 2 + ImageColorizer Original
Static (Mip-NeRF-360) 8 mins 31 mins 50 mins 42 mins 16 h
Dynamic (DyNeRF) 8 mins 38 mins 54 mins 47 mins
Generalization on dissimilar scenes. As discussed in Section 4.5 and illustrated in Fig. 7, our personalized colorizer can be transferred to similar scenes without retraining, which can reduce the optimization burden for large-scale similar scenes. Here, we further evaluate the performance of the proposed method on more challenging, substantially different scenes. Specifically, we train the colorizer on the “Garden” scene from Mip-NeRF 360 and transfer it to the “Room” scene. As shown in Fig. 17, our method achieves suboptimal performance when the target scene is very different from the training scene. For example, some sofa regions inherit green tones from the source scene, and the overall color richness is reduced. Nevertheless, the colorizer still produces overall consistent and stable results and, in particular, performs reasonably well for elements that share similar semantic properties, such as the vegetation outside the window. In summary, the personalized colorizer demonstrates potential generalization to similar scenes, reducing the need for retraining across many visually similar scenes in large-scale datasets. However, we still advocate scene-specific fine-tuning to achieve optimal 3D colorization results, which remains our recommended practice in this work.
Figure 17: Visualization of the generalization performance on dissimilar scene.
@@
Image and video colorizers vs. per-scene colorizer. We evaluate the effectiveness of our per-scene colorizer design by comparing it with two representative baselines: the image colorization model DDColor Kang et al. (2023) and the video colorization method ColorMNet Yang et al. (2024a), as illustrated in Fig. 18. It can be observed that DDColor suffers from severe cross-view color inconsistencies due to its per-image one-to-many color mapping nature. Treating multi-view images as a video sequence and applying video colorization similarly fails to resolve this issue, since the viewpoint changes inherent in 3D capture far exceed the temporal variations commonly seen in video data. In contrast, our personalized colorizer learns a scene-specific, one-to-one color mapping from the reference view and leverages its inherent inductive bias to consistently project colors to corresponding content in novel views and time steps. This design effectively addresses the challenge of color consistency in both static and dynamic 3D scenes.
@@
Effort of 3D representations. We further investigate how different 3D representations impact colorization performance. As shown in Tab.8, compared to 3DGS, combining NeRFMildenhall et al. (2021) or NeRFPlayer Song et al. (2023) with image colorizers yields suboptimal consistency, reflected by an increase of 0.023 in ME. This is mainly because the explicit modeling of 3DGS gives it an inherent resistance and geometric priors, which can naturally alleviate color inconsistency to some extent. In addition, we replace the Lab Gaussian representation in Color3D with NeRF and NeRFPlayer, respectively. The results underscore the effectiveness of our proposed Lab Gaussian representation while also demonstrating that the Color3D framework can be adapted to various 3D reconstruction backbones with reasonable performance.
@@
Method Language-Guided Colorization
@@
NeRF+ImageColorizer 104.85 0.6153 0.225 0.046
3DGS+ImageColorizer 112.33 0.6148 0.202 0.042
Color3D-NeRF 76.68 0.6213 0.065 0.014
Color3D (Ours) 68.23 0.6246 0.058 0.012
@@
NeRFPlayer+ImageColorizer 86.77 0.6148 0.133 0.027
4DGS+ImageColorizer 89.39 0.6159 0.124 0.025
Color3D-NeRFPlayer 63.29 0.6243 0.046 0.009
Color3D (Ours) 58.62 0.6271 0.041 0.008
@@
User study. To further assess the perceptual quality of the generated results, we conduct a comprehensive user study, acknowledging that subjective human perception remains a critical benchmark in the field of colorization. Specifically, we design 5 evaluation questions covering multiple aspects, including color richness, consistency, aesthetic preference, image quality, and the fidelity of alignment between the synthesized views and the provided conditions (text, exemplars). Participants are asked to rate each sample on a scale of 0 to 5 for each aspect. For a fair comparison, we recruit 30 participants, each evaluating 10 sets of comparisons across different methods. Aggregated results are shown in Fig. 21. The naive combination of 3DGS with off-the-shelf image colorizers yields reasonably rich colors but suffers from significant inconsistency and reduced visual quality. While ColorNeRF demonstrates better consistency, it often lacks color diversity and visual appeal. In contrast, our method consistently achieves higher user ratings across all aspects, highlighting its advantage in terms of perceptual quality and user preference.
@@
A.6 Limitations and Future Work